1) A bag contains 10 purple marble, 15 yellow marbles and 5 orange marbles . You are going to remove one marble from the bag without looking. What is the probability that , you will choose a yellow marble ?

2)Twenty -three markers are put in a straight line 14cm apart from each other. How far is it from the first marker to the last?
Answer #1

1)

Probability (P) = Favourable Outcomes(F) / Total Outcomes (T)

Total Outcomes = 30C1 = 30

Favourable Outcomes = 15C1 =15

Hence, P = 15/30 = 0.5 or 50%

2)

Per unit distance = 14cm

n = 23

Total Distance = (n-1)*p = 22*14 = 308 cms